# Language-Integrated Recursive Queries

### v1

@@ 4.4. Linear Recursion
We approach the problem of encoding linear recursion in the type system with a linear type with respect to recursive references. To accomplish this, we split the problem into two sub-constraints: an affine check (every relation is never used more than once) and a relevant check (all relations are used in at least one recursive definition). The affine check is per-relation, while the relevant check is for all relations, as not every relation needs to use every other relation within a single stratum. The extensions to the type system to restrict queries to linear recursion are shown in Figure 9.
A type parameter of type is added to to model the dependencies of each query. Uniqueness of references is enforced by using the argument position within the fix function: for a recursive function , will be a tuple containing the constant integer type , and will contain a tuple with the constant integer type .
Some database backends, for example DuckDB, allow recursive queries nested within outer queries to return columns from the outer query (which can itself be recursive, e.g. fix within a fix). However, this would break linearity, so must be able to differentiate between references per call to fix. Each reference is tagged with a singleton type that is unique for each fix invocation. References are considered equal only if both the and the constant integer type are the same. This ensures that if fix is called within the scope of references from another fix function, the inner fix cannot return terms derived from the outer function’s parameters. For example:
For the affine check, the dependencies of each recursive definition must contain no duplicates. For the relevant check, all parameters must appear at least once in all . This is implemented by using the set of constant integer types from 1 to the number of arguments to fix decorated with to check that all elements are present at least once in all the s () and never more than once in each ().
@@
If then holds the linearity property if it holds both the relevant and affine properties. The relevant property holds if and the affine property if .
4.5. Set semantic Recursion
| Syntax | |||||
@@ 4.5. Set semantic Recursion
To restrict fix to allow only set semantics within recursive queries, we track the semantics of each operation in a type parameter, similarly to how we tracked monotonicity. We update so that all relation-level operations reflect their semantics. The required changes to the syntax and typing rules are presented in Figure 10. For example:
@@
A function such that holds the set-semantic property if .
4.6. Constructor-Freedom
@@
A function such that holds the constructor-freedom property if and is of type RestrictedExpr.
@@ 4.7. Nesting in
The problem of deeply nested datatypes and query avalanches is well covered by previous work in NRC and T-LINQ. The unnesting and normalization approach taken by T-LINQ is applicable to non-recursive queries, or solely within the bodies of recursive queries, but it is not applicable between non-recursive and recursive subqueries because the fixpoint introduces a strict evaluation boundary—often referred to as a ‘pipeline breaker’, which prevents the same normalization approach. Therefore, we have chosen to simplify the syntax of to restrict nested column types, although in our implementation (discussed in Section 6) nested datatypes are supported.
@@
The power of language-integrated query comes from the tight embedding of the DSL with the general-purpose host language. The DSL environment represents staged computation: all functions in the DSL syntax serve to construct a type-level AST but do not actually execute any database operations. To generate queries that are run on a database and returned to an application written in a general-purpose language, we need to extend to be an embedded DSL in a host language. In Figure 11, based on the precedent set by T-LINQ, we extend with a second type environment for host terms, , and update our DSL typing rules accordingly. Judgment states that host term has type in type environment , and judgment states that quoted term has type in host type environment and DSL type environment .
We reuse the syntax of types and terms in Figure 5 for the host language, with a few key additions. DSL types are represented in the host language by wrapping them in Expr types, so if a DSL expression represents a database row of type then the type of this expression in the host language will be . Translation functions from host terms to DSL terms are added: toExpr converts base host-language types to Expr of base types and toRow converts Named-Tuples of Expr to Exprs of Named-Tuple. Finally, an execution function run is added to execute the query expressed in the DSL and returns a List of the query result type to the application. TO-EXPR, TO-ROW, RUN-QUERY, and RUN-AGG rules define the types for toExpr, toRow, and run. The typing rules for in Figure 11 combine all the restrictions presented in Figures 8, 9, 10 with the addition of rules to handle interactions between the DSL and host language. The typing rule for the fully-restricted host language fix is shown in Figure LABEL:fig:fix-impl.
@@ 6. Implementation
In this section, we describe the implementation of TyQL, our type-safe embedded query library based on . Figure LABEL:fig:fix-impl shows the implementation of fix in TyQL with all restrictions applied and the typing rule from with all restrictions applied, renamed to match the variable names in TyQL for readability. We include the side-by-side comparison to emphasize that our implementation and our theory are closely knit. The architecture of TyQL is summarized in Figure 12.
@@ 6.2.2. Projection
The method that models projection is defined in Figure 14. Type-level pattern matching with Match Types (Blanvillain et al., 2022) is used to lift tuples-of-expressions to expressions-of-tuples (lines 1-3) and to enforce that all values of the tuple are of type Expr (lines 4-5). The IsTupleOfExpr constraint first converts the Named-Tuple into a regular Tuple with NamedTuple.DropNames, then takes the Union type of all values in the tuple with Tuple.Union, and finally constrains the resulting Union type to be a subtype of a non-scalar Expr with <:<. The toRow method is defined (lines 6-7) on Named-Tuples of type Expr and constructs an instance of the Project AST node. This method is used as an implicit conversion between Named-Tuple-of-Expr and Project[A] (line 8). TyQL reifies the constant string keys in Named-Tuples using type tags (ResultTag) so that the generated queries are more readable, for example t.map(r => (maxCount = max(r.count)) generates the SQL query SELECT MAX(count) as maxCount FROM t.
@@ 6.2.5. Recursive Constraints
Query syntax trees are built and verified by the host language compiler, and constraints are enforced with type classes and implicit evidence. As shown in Figure LABEL:fig:fix-impl, each of the implicit constraints in the definition of fix in TyQL have a corresponding premise in (with the exception of line #7, which is essentially a hint to the compiler to understand the structure of the dependency tuple, DT). TyQL uses the annotation @implicitNotFound to customize error messages, so that the user can see which constraint failed: i.e., if their recursive query was affine but not relevant. For systems that do not support mutual recursion, we do not have to include this constraint in the type system because we can simply limit the number of arguments to the fix function to a single relation, which prevents the construction of mutually recursive queries. Unique query references are implemented using constant integer types and unique types are implemented using anonymous classes and value types.
Figure LABEL:fig:fix-impl shows the fully restricted fix, but the restrictions can be made configurable: for example, if a user wants a fix function with all restrictions except set-semantics, they simply call val bagFixFn = fix((ConstructorFree(), Monotone(), BagResult(), Linear())). This is implemented by adding additional type parameters for each restriction [RCF <: ConstructorFreedom, RM <: MonotoneRestriction, RC <: ResultCategory, RL <: LinearRestriction] to line 2 and an additional argument to line 3: (options: (RCF, RM, RC, RL)). The type arguments are forwarded to the match types ToRestrictedQueryRef and ToRestrictedQuery in lines 4 and 8 that will evaluate to different types depending on those arguments.
@@ 7. Evaluation
| ⬇ // TyQL: val base = Table[(dst:Int, cst:Int)]("base") val edges = Table[(src:Int, dst:Int, cst:Int)]("edge") base.fix(sp => edges.flatMap(edge => sp.filter(s => s.dst == edge.src) .map(s => (dst = edge.dst, cst = s.cst + edge.cst)) .distinct) .aggregate(s => (dst = s.dst, cst = min(s.cst))) .groupBy(s => (dst = s._1.dst)) | ⬇ -- SQL: CREATE TABLE edge (src INT, dst INT, cst INT); CREATE TABLE base (dst INT, cst INT); WITH RECURSIVE path AS ( (SELECT * FROM base) UNION (SELECT edge.dst, path.cst + edge.cst FROM edge, path WHERE path.dst = edge.src)) SELECT path.dst, MIN(path.cst) FROM path GROUP BY path.dst |
SSSP Query.
| Query | Description | Linear | Mono-tone | Set | Mut-ual | CF |
| Even-Odd | ◼ Mutual recurive program generates even/odd numbers. | ✓ | ✓ | ✓ | ✓ | ✗ |
| CSPA | ⚫ Graspan’s Context Sensitive Pointer Analysis, static analysis for identifying aliases. (Wang et al., 2017) | ✗ | ✓ | ✓ | ✓ | ✓ |
| Company Control (CC) | ▲ Calculates the complex controlling relationships between companies. (Mumick et al., 1990) | ✓ | ✗ | ✓ | ✓ | ✗ |
| PointsToCount (PTC) | ⚫ Find the count of objects that a variable of a given name may point to (Abeysinghe et al., 2024) (Java variant). | ✗ | • | ✓ | ✓ | ✓ |
| Chain Of Trust (COT) | ◼ Security query where two entities trust each other if they are direct friends or if their friends trust each other. | ✓ | ✓ | ✗ | ✓ | ✓ |
| Java Points To (JPT) | ⚫ Field-sensitive subset-based points to analysis for an object-oriented language. (Madsen et al., 2016) | ✗ | ✓ | ✗ | ✓ | ✓ |
| Party | ★ Mutually recursive social media graph algorithm.(Wang et al., 2020) | ✓ | ✗ | ✗ | ✓ | ✗ |
| CBA | ◼ Constraint-based analysis query. (souffle, 2019) | ✗ | • | ✗ | ✓ | ✓ |
| Single Source Shortest Path (SSSP) | ★ Computes the shortest path from a given source node to all other nodes in a weighted graph. (Wu et al., 2022) | ✓ | • | ✓ | ✗ | ✗ |
| Ancestry | ▲ Find descendants of a person who are of the same generation. (Bancilhon and Ramakrishnan, 1986) | ✓ | ✓ | ✓ | ✗ | ✗ |
| Andersen’s Points To (APT) | ⚫ Context-insensitive, flow-insensitive interprocedural pointer analysis. (Fan et al., 2019a) | ✗ | ✓ | ✓ | ✗ | ✓ |
| All Pairs Shortest Path (APSP) | ★ Compute the shortest paths between all pairs of nodes in a weighted graph. (Wu et al., 2022) | ✗ | ✗ | ✓ | ✗ | ✗ |
| Graphalytics (TC) | ★ Directed cyclic graph reachability query. Uses list data structure to check for cycles. (Angles et al., 2020) | ✓ | ✓ | ✗ | ✗ | ✗ |
| Bill of Materials (BOM) | ▲ Business query for how many days it would take to deliver a product made of subparts. (ibm, 2021) | ✓ | • | ✗ | ✗ | ✓ |
| Orbits | ◼ Orbits of cosmological objects. (souffle, 2019) | ✗ | • | ✗ | ✗ | ✓ |
| Data Flow | ◼ Models control flow through read/write/jump instructions. (souffle, 2019) | ✗ | ✓ | ✗ | ✗ | ✓ |
Linear, Set, Mutual, CF (constructor-free): ✓ query has the property.
Monotone: ✓ no aggregation • stratified aggregation ✗ unstratified aggregation.
In this section, we evaluate TyQL with respect to real-world use cases and how effectively TyQL can constrain recursive queries without compromising on usability. To illustrate the range of real-world queries, we introduce a recursive query benchmark (RQB) of 16 queries taken from open source code repositories and recent publications from various query domains, including business management, program analysis, graph queries, and other classic fixed-point problems. We give a survey of modern relational database systems using our benchmark, identifying the ways each query can go wrong and which RDBMS supports each query property or combination of properties.
@@ 7.1. Recursive Query Benchmark and Survey of RDBMS
In this section, we present a recursive query benchmark (RQB) comprising 16 representative queries across diverse domains such as business management, program analysis, graph queries, and classic fixed-point problems and explore how each query behaves on different RDBMS.
@@
The goal of our benchmark is to simulate a broad range of recursive queries. We have identified five discrete properties of recursive queries that have the greatest impact on behavior: (1) linearity, (2) monotonicity, (3) set semantics, (4) mutual recursion, and (5) constructor-freedom. All queries are range-restricted. We have selected 16 queries to represent classes of queries for each combination of properties 1-4. Since constructor-freedom applies independently to any query, we included a mix of queries with this property. Identifying the frequency of each query across real-world applications remains future work. All queries run and terminate on at least one of the evaluated database systems. For the monotonicity property, we have selected a mix of queries without aggregation or negation, with stratified-aggregation, and with unstratified aggregation. Table 2 illustrates the benchmark property matrix. As an example, we illustrate the Single Source Shortest Path query expressed in TyQL and in SQL in Figure 15. The full set of queries in the recursive benchmark expressed in both TyQL and the generated SQL are included in the appendix.
| DuckDB | Postgres | SQLite | MariaDB | What restriction is violated? | |
| Even-Odd | ✓ | ✗ | ✗ | ◗ | mutual, cf |
| CSPA | ◗ | ✗ | ✗ | ✗ | linear, mutual |
| CC | ◗ | ✗ | ✗ | ✗ | monotone, mutual, cf |
| PTC | ◗ | ✗ | ✗ | ✗ | linear, mutual |
| COT | ✓ ❘ ▼✳ | ✗ | ✗ | ✗ | set, mutual |
| JPT | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ | linear, set, mutual |
| Party | ◗ | ✗ | ✗ | ▼ | monotone, set, mutual, cf |
| CBA | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ | linear, set, mutual |
| SSSP | ✓ ❘ ▼✳ | ✓ ❘ ▼ | ✓ ❘ ▼ | ✓ | cf |
| Ancestry | ✓ | ✓ | ✓ | ✓ | cf |
| APT | ◗ | ◗ | ✗ | ◗ | linear |
| APSP | ◗ | ◗ | ✗ | ✓ | linear, monotone, cf |
| TC | ✓ | ✓ | ✓ | ✓ | set, cf |
| BOM | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ | set |
| Orbits | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ | linear, set |
| Data Flow | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ | linear, set |
@@
As shown in Section 2.2, support for recursive queries varies widely across RDBMS. To get a sense of what classes of queries each system supports, we ran the 16 queries with cyclic input data and with acyclic input data. The results of the RQB on four RDBMS are presented in Table 3. We represent queries that terminated with the full result with ✓, database error with ✗, incomplete results with ◗, and nontermination with ▼. If the query exhibited different behavior based on if the input data was cyclic or not, for example if the query terminates correctly with acylic data but not cyclic data, then the query is represented as ✓ ❘ ▼ with the acyclic data on the left and cyclic data on the right. We consider “incomplete” results to be data that is missing results that would have been present using a non-SQL version of each algorithm, either via general-purpose imperative programs or Datalog programs. The • and ▼ classification indicates that it is possible to find an input dataset that returns incomplete results or does not terminate. The databases used are DuckDB v1.1, Postgres v15, SQLite v3.39, and MariaDB v11.5.2 with the configuration ---skip-standard-compliant-cte.

### v2

@@ 5.5. P4: Linear Recursion
We approach the problem of encoding linear recursion in the type system with a linear type with respect to recursive relations. To accomplish this, we split the problem into two sub-constraints: an affine check (every recursive relation is never used more than once to define a recursive relation) and a relevant check (all recursive relations are used in at least one recursive definition). The affine check is per-relation, while the relevant check is for all relations defined within a single fix call, as not every relation needs to use every other relation within a single stratum. The extensions to the type system to restrict queries to linear recursion are shown in Figure 9.
A type parameter of type is added to to model the dependencies of each query. We use to indicate multiset union, so duplicates are maintained, and equivalence using does not take order into account. Uniqueness of references is enforced by using the argument position within the fix function: for a recursive function , will be a tuple containing the constant integer type , and will contain a tuple with the constant integer type . Some database backends, for example DuckDB, allow recursive queries nested within outer queries to return columns from the outer query (which can itself be recursive, e.g. fix within a fix). However, this would break linearity, so must be able to differentiate between references per call to fix. Each reference is tagged with a singleton type that is unique for each fix invocation. References are considered equal only if both the and the constant integer type are the same. This ensures that if fix is called within the scope of references from another fix function, the inner fix cannot return terms derived from the outer function’s parameters. Multi-relation operations like flatMap or union collate recursive references in the result relation. For example:
For the affine check, the dependencies of each recursive definition must contain no duplicates. This check is implemented by taking the length (indicated with ) of and requiring that is the same as the length of the union of : . For the relevant check, all parameters must appear at least once in all . This is implemented by using the set of constant integer types from 1 to the number of arguments to fix decorated with to check that all elements are present at least once in all the s: .
@@
If then holds the linearity property P4 if it holds both the relevant and affine properties. The relevant property holds if and the affine property if .
5.6. P5: Set-semantic Recursion
| Syntax | ||||||||
@@ 5.6. P5: Set-semantic Recursion
To restrict fix to allow only set semantics within recursive queries, we track the semantics of each query-level operation in a type parameter, similarly to how we tracked monotonicity in column-level operations. We update so that all query-level operations reflect their semantics. The required changes to the syntax and typing rules are presented in Figure 10. For example:
@@
A function such that holds the set-semantic property P5 if .
5.7. P6: Constructor-Freedom
@@
A function such that holds the constructor-freedom property P6 if and is of type RestrictedExpr.
@@ 5.8. Nesting in
The problem of deeply nested datatypes and query avalanches is well covered by previous work in NRC and T-LINQ. The unnesting and normalization approach taken by T-LINQ is applicable to non-recursive queries, or solely within the bodies of recursive queries, but it is not applicable between non-recursive and recursive subqueries because the fixpoint introduces a strict evaluation boundary - often referred to as a ‘pipeline breaker’, which prevents the same normalization approach. Therefore, we have chosen to simplify the syntax of to restrict nested column types, although in our implementation (discussed in Section 7) nested datatypes are supported.
@@
The power of language-integrated query comes from the tight embedding of the DSL with the general-purpose host language. The DSL environment represents staged computation: all functions in the DSL syntax serve to construct a type-level AST but do not actually execute any database operations. To generate queries that are run on a database and returned to an application written in a general-purpose language, we need to extend to be an embedded DSL in a host language.
In Figure 11, based on the precedent set by T-LINQ, we extend with a second type environment for host terms, , and update our DSL typing rules accordingly. Judgment states that host term has type in type environment , and judgment states that quoted term has type in host type environment and DSL type environment .
We reuse the syntax of types and terms in Figure 6 for the host language, with a few key additions. Lambda application, i.e., , is added to the host-language syntax with the typing rule APP so users can define and use arbitrary functions. DSL types are represented in the host language by wrapping them in Expr types, so if a DSL expression represents a database row of type then the type of this expression in the host language will be . Translation functions from host terms to DSL terms are added: toExpr converts base host-language types to Expr of base types and toRow converts Named-Tuples of Expr to Exprs of Named-Tuple. Finally, an execution function run is added to execute the query expressed in the DSL and returns a List of the query result type to the application. TO-EXPR, TO-ROW, RUN-QUERY, and RUN-AGG rules define the types for toExpr, toRow, and run. The typing rules for in Figure 11 combine all the restrictions presented in Figures 8, 9, 10 with the addition of rules to handle interactions between the DSL and host language. The typing rule for the fully-restricted host language fix is shown in Figure LABEL:fig:fix-impl (to restrict mutual recursion, we limit the meta-syntax to ).
@@ 7. Implementation
In this section, we describe the implementation of TyQL, our type-safe embedded query library based on . Figure LABEL:fig:fix-impl shows the implementation of fix in TyQL with all restrictions applied except mutual recursion and the typing rule from , renamed to match the variable names in TyQL for readability. We include the side-by-side comparison to emphasize that our implementation and our theory are closely knit. The architecture of TyQL is summarized in Figure 12.
@@ 7.2.2. Projection
The method that models projection is defined in Figure 14. Type-level pattern matching with Match Types (Blanvillain et al., 2022) is used to lift tuples-of-expressions to expressions-of-tuples (lines 1-3) and to enforce that all values of the tuple are of type Expr (lines 4-5). The IsTupleOfExpr constraint first converts the Named-Tuple into a regular Tuple with NamedTuple.DropNames, then takes the Union type of all values in the tuple with Tuple.Union, and finally constrains the resulting Union type to be a subtype of a non-scalar Expr with <:<. The toRow method is defined (lines 6-7) on Named-Tuples of type Expr and constructs an instance of the Project AST node. This method is used as an implicit conversion between Named-Tuple-of-Expr and Project[A] (line 8). TyQL reifies the constant string keys in Named-Tuples using type tags (ResultTag) so that the generated queries are more readable, for example r.map(r => (maxCount = max(r.count)) generates the SQL query SELECT MAX(count) as maxCount FROM R.
@@ 7.2.5. Recursive Constraints
Query syntax trees are built and verified by the host language compiler, and constraints are enforced with type classes and implicit evidence. As shown in Figure LABEL:fig:fix-impl, each of the implicit constraints in the definition of fix in TyQL have a corresponding premise in (with the exception of line #7, which is essentially a hint to the compiler to understand the structure of the dependency tuple, DT). TyQL uses the annotation @implicitNotFound to customize error messages, so that the user can see which constraint failed: i.e., if their recursive query was affine but not relevant. To restrict mutual recursion in Figure LABEL:fig:fix-impl, we could tweak fix such that QT, DT, and RQT must be subtypes of Tuple1 (tuples of length 1). However, we prefer to provide a more erognomic implementation where fix is a method of Query, so that users can simply write edges.fix(e -> ...). Unique query references are implemented using constant integer types and unique types are implemented using anonymous classes and value types.
The restrictions of P1-P6 can be made configurable: for example, if a user wants a fix function with all restrictions except set-semantics, they call val bagFixFn = fix((ConstructorFree(), Monotone(), BagResult(), Linear())). This is implemented by adding additional type parameters for each restriction [RCF <: ConstructorFreedom, RM <: MonotoneRestriction, RC <: ResultCategory, RL <: LinearRestriction] to line 2 and an additional argument to line 3: (options: (RCF, RM, RC, RL)). The type arguments are forwarded to the match types ToRestrictedQueryRef and ToRestrictedQuery in lines 4 and 8 that will evaluate to different types depending on those arguments.
@@ 7.3. Query Generation
With respect to query construction, TyQL doesn’t introduce any additional overheads from JVM boxing beyond what Scala does when working with generic ADTs or case classes. With respect to results sent back from the database, all boxing is handled by the JDBC driver regardless of if the query was constructed with strings or with TyQL.
@@ 8. Evaluation
In this section, we evaluate TyQL with respect to real-world use cases and how effectively TyQL can constrain recursive queries without compromising on usability. To illustrate the range of real-world queries, we introduce a recursive query benchmark (RQB) of 16 queries taken from open source code repositories and recent publications from various query domains, including business management, program analysis, graph queries, and other classic fixed-point problems. We give a survey of modern relational database systems using our benchmark, identifying the ways each query can go wrong and which RDBMS supports each query property or combination of P1-P6.
@@ 8.1. Recursive Query Benchmark and Survey of RDBMS
| Query | Description | P2 | P3 | P4 | P5 | P6 |
| Even-Odd | ◼ Mutually recursive program generates even/odd numbers. | ✓ | ✓ | ✓ | ✓ | ✗ |
| CSPA | ⚫ Graspan’s Context Sensitive Pointer Analysis, static analysis for identifying aliases. (Wang et al., 2017) | ✓ | ✓ | ✗ | ✓ | ✓ |
| Company Control (CC) | ▲ Calculates the complex controlling relationships between companies. (Mumick et al., 1990) | ✗ | ✓ | ✓ | ✓ | ✗ |
| PointsToCount (PTC) | ⚫ Find the count of objects that a variable of a given name may point to (Abeysinghe et al., 2024) (Java variant). | • | ✓ | ✗ | ✓ | ✓ |
| Chain Of Trust (COT) | ◼ Security query where two entities trust each other if they are direct friends or if their friends trust each other. | ✓ | ✓ | ✓ | ✗ | ✓ |
| Java Points To (JPT) | ⚫ Field-sensitive subset-based points to analysis for an object-oriented language. (Madsen et al., 2016) | ✓ | ✓ | ✗ | ✗ | ✓ |
| Party | ★ Mutually recursive social media graph algorithm.(Wang et al., 2020) | ✗ | ✓ | ✓ | ✗ | ✗ |
| CBA | ◼ Constraint-based analysis query. (souffle, 2019) | • | ✓ | ✗ | ✗ | ✓ |
| Single Source Shortest Path (SSSP) | ★ Computes the shortest path from a given source node to all other nodes in a weighted graph. (Wu et al., 2022) | • | ✗ | ✓ | ✓ | ✗ |
| Ancestry | ▲ Find descendants of a person who are of the same generation. (Bancilhon and Ramakrishnan, 1986) | ✓ | ✗ | ✓ | ✓ | ✗ |
| Andersen’s Points To (APT) | ⚫ Context-insensitive, flow-insensitive interprocedural pointer analysis. (Fan et al., 2019) | ✓ | ✗ | ✗ | ✓ | ✓ |
| All Pairs Shortest Path (APSP) | ★ Compute the shortest paths between all pairs of nodes in a weighted graph. (Wu et al., 2022) | ✗ | ✗ | ✗ | ✓ | ✗ |
| Graphalytics (TC) | ★ Directed cyclic graph reachability query. Uses list data structure to check for cycles. (Angles et al., 2020) | ✓ | ✗ | ✓ | ✗ | ✗ |
| Bill of Materials (BOM) [Stratified] | ▲ Business query for days to deliver a product made of subparts with stratified aggregation. (ibm, 2021) | • | ✗ | ✓ | ✗ | ✓ |
| Orbits | ◼ Orbits of cosmological objects. (souffle, 2019) | • | ✗ | ✗ | ✗ | ✓ |
| Data Flow | ◼ Models control flow through read/write/jump instructions. (souffle, 2019) | ✓ | ✗ | ✗ | ✗ | ✓ |
Benchmark domain: ◼ Datalog ⚫ Program analysis ▲ Recursive SQL ★ Graph.
P2 (monotone): ✓ no aggregation • stratified aggregation ✗ unstratified aggregation.
P3 (mutually-recursive), P4 (linear), P5 (set-semantic), P6 (constructor-free): ✓ query has the property.
@@
In this section, we present a recursive query benchmark (RQB) comprising 16 queries across diverse domains such as business management, program analysis, graph queries, and classic fixed-point problems and show how each query behaves on different RDBMS.
@@
The goal of our benchmark is to simulate a broad range of recursive queries. We have selected 16 queries to represent classes of queries for each combination of properties (except P1 as all queries are range-restricted). We excluded trivially safe queries (those violating none of P1–P6) because they pose no risk of causing behaviors B1-B3. Identifying the frequency of each query across real-world applications remains future work. All queries run and terminate on at least one of the evaluated RDBMS. For the monotonicity property, we have selected a mix of queries without aggregation or negation, with stratified-aggregation, and with unstratified aggregation. Table 2 illustrates the benchmark property matrix. As an example, we illustrate the Single Source Shortest Path query expressed in TyQL and in SQL in Figure 15. The full set of queries in the recursive benchmark expressed in both TyQL and SQL are included in the appendix.
| ⬇ // TyQL: val base = Table[(dst:Int, cst:Int)]("base") val edges = Table[(src:Int, dst:Int, cst:Int)]("edge") base.fix(path => edges.flatMap(edge => path.filter(s => s.dst edge.src) .map(s => (dst = edge.dst, cst = s.cst + edge.cst)) .distinct) .aggregate(s => (dst = s.dst, cst = min(s.cst))) .groupBy(s => (dst = s._1.dst)) | ⬇ -- SQL: CREATE TABLE Edges (src INT, dst INT, cst INT); CREATE TABLE Base (dst INT, cst INT); WITH RECURSIVE Path AS ( (SELECT * FROM Base) UNION (SELECT Edges.dst, Path.cst + Edges.cst FROM Edges, Path WHERE Path.dst = Edges.src)) SELECT Path.dst, MIN(Path.cst) FROM Path GROUP BY Path.dst |
SSSP Query.
| Benchmark | Database Behavior | ||||
| Query | Violated Properties | DuckDB | Postgres | SQLite | MariaDB |
| Even-Odd | P3, P6 | ✓ | ✗ | ✗ | ◗ |
| CSPA | P3, P4 | ◗ | ✗ | ✗ | ✗ |
| CC | P2, P3, P6 | ◗ | ✗ | ✗ | ✗ |
| PTC | P3, P4 | ◗ | ✗ | ✗ | ✗ |
| COT | P3, P5 | ✓ ❘ ▼✳ | ✗ | ✗ | ✗ |
| JPT | P3, P4, P5 | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ |
| Party | P2, P3, P5, P6 | ◗ | ✗ | ✗ | ▼ |
| CBA | P3, P4, P5 | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ |
| SSSP | P6 | ✓ ❘ ▼✳ | ✓ ❘ ▼ | ✓ ❘ ▼ | ✓ |
| Ancestry | P6 | ✓ | ✓ | ✓ | ✓ |
| APT | P4 | ◗ | ◗ | ✗ | ◗ |
| APSP | P2, P4, P6 | ◗ | ◗ | ✗ | ✓ |
| TC | P5, P6 | ✓ | ✓ | ✓ | ✓ |
| BOM | P5 | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ |
| Orbits | P4, P5 | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ |
| Data Flow | P4, P5 | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ |
@@
As shown in Section 2, support for recursive queries varies widely across RDBMS. To get a sense of what classes of queries each system supports, we ran the 16 queries with cyclic input data and with acyclic input data. The results of the RQB on four RDBMS are presented in Table 3. We represent queries that terminated with the full result with ✓, database error (B1) with ✗, incomplete results (B2) with ◗, and nontermination (B3) with ▼. If the query exhibited different behavior based on if the input data was cyclic or not, for example if the query terminates correctly with acylic data but not cyclic data, then the query is represented as ✓ ❘ ▼ with the acyclic data on the left and cyclic data on the right. We consider “incomplete” results to be data that is missing results present using a non-SQL version of each algorithm, either via imperative programs or Datalog programs. The • and ▼ classification indicates that it is possible to find an input dataset that returns incomplete results or does not terminate. The databases used are DuckDB v1.1, Postgres v15, SQLite v3.39, and MariaDB v11.5.2 with the configuration ---skip-standard-compliant-cte.